OpenStudy (anonymous):

If g(x) = 2x- x +5 , then g (2) =

OpenStudy (anonymous):

The answer choices are a.2 b.6 c.7 d.9 e.19

OpenStudy (anonymous):

g(2)=2(2)-(2)+5=4-2+5=7

OpenStudy (anonymous):

since g(x) became g(2) that means x=2 so you just have to put value of 2 into the place of x

OpenStudy (anonymous):

Thank you!!